# A Damage-Driven Model for Duchenne Muscular Dystrophy: Early-Stage Dynamics and Invasion Thresholds

### 31 Mar 2026 v2

@@ Abstract
We introduce a spatially extended mathematical model for Duchenne muscular dystrophy based on a damage-driven paradigm, in which immune recruitment is triggered by tissue injury. The model is formulated as a reaction–diffusion–chemotaxis system describing the interaction between healthy tissue, damaged fibers, immune cells and inflammatory signals. We establish the global well-posedness of the system and investigate the early-stage dynamics through linearization around the healthy equilibrium. Our analysis shows that diffusion does not induce Turing instabilities, so that spatial heterogeneity cannot arise from diffusion-driven mechanisms. Instead, disease progression occurs through invasion processes. We derive explicit conditions for the onset of invasion, interpreted as an effective damage reproduction threshold and characterize the minimal propagation speed of pathological fronts, showing that the dynamics is governed by a pulled-front mechanism. Numerical simulations support the analytical results and confirm the transition between decay and invasion. These results provide a mathematical framework for early-stage disease progression and indicate that spatial patterns arise from the expansion of localized damage rather than from intrinsic pattern-forming mechanisms.
@@ 3.2.2 Pathological equilibria
4 Well-posedness of the reaction–diffusion–chemotaxis system
In this section, we establish the global well-posedness of the reaction–diffusion–chemotaxis system (2.3). In particular, we prove the existence, uniqueness and boundedness of solutions, together with the preservation of a biologically admissible domain, ensuring that all components remain non-negative and that the saturation constraint $h+d\leq 1$ is satisfied for all times. These results hold independently of the specific dynamical regime and provide the analytical framework required for the subsequent investigation of the system dynamics.
The analysis relies on a Leray–Schauder fixed point argument, combined with a priori estimates and parabolic regularity techniques. These results provide the mathematical framework required for the subsequent investigation of the dynamical properties of the model.
@@ 4 Well-posedness of the reaction–diffusion–chemotaxis system
$\left\{\begin{array}[]{ll}\displaystyle\partial_{t}h=\Delta h+(\sigma+\rho d)(1-h-d)-h-A(m,c)h\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}d=D_{d}\Delta d+A(m,c)h-\delta d-\rho d(1-h-d)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(d-m)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}c=D_{c}\Delta c+rd-\mu c,&(t,x)\in\Omega_{T},\end{array}\right$. (4.1)
with
$\chi(m,c)=\chi_{0}\dfrac{m\,c}{\kappa+c},\quad A(m,c)=\alpha\dfrac{c+c_{\epsilon}}{1+c}\,m\,$, (4.2)
and we denote by $\Omega_{T}=(0,T)\times\Omega$, where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ $(n\in\mathbb{N},\ n\geq 1)$ with smooth, specified later, boundary $\partial\Omega$. We denote by $\mathbf{n}=\mathbf{n}(s)$ the outward normal to $\partial\Omega$ at a point $s$ of the boundary, and impose the following Neumann boundary condition on $\partial\Omega$:
$\begin{cases}\mathbf{n}(s)\cdot\bm{\nabla}h(t,s)=0\,,&\qquad(t,s)\in(0,T)\times\partial\Omega\,,\\ \mathbf{n}(s)\cdot\bm{\nabla}d(t,s)=0\,,&\qquad(t,s)\in(0,T)\times\partial\Omega\,,\\ \mathbf{n}(s)\cdot\bm{\nabla}m(t,s)=0\,,&\qquad(t,s)\in(0,T)\times\partial\Omega\,,\\ \mathbf{n}(s)\cdot\bm{\nabla}c(t,s)=0\,,&\qquad(t,s)\in(0,T)\times\partial\Omega\,.\end{cases}$ (4.3)
@@
[…]
and we suppose that the following condition holds
@@
Moreover, we recall that all system parameters are positive.
We introduce the functional framework used throughout the section (see [14, 21]).
@@
We denote by $W^{r,p}(\Omega)$ the Sobolev spaces of order $r$ with integrability $p$. For the parabolic setting, we consider the space
@@
endowed with the norm
@@
We also denote by $C^{k}(\overline{\Omega})$ the space of functions with continuous derivatives up to order $k$, and by $C^{0}(\Omega_{T})$ the space of continuous functions in $\Omega_{T}$.
Finally, we introduce the space $C^{0,1}=C^{0,1}([0,T]\times\overline{\Omega})$ of functions that are continuous together with their spatial gradients, endowed with the norm
@@
$\partial_{t}(h,d,m,c)\,,\,\,\Delta(h,d,m,c)\,\in[L^{\tilde{p}}\left(\Omega_{T}\right)]^{4}\,\,,\,\,\nabla\cdot\left(\chi(m,c)\nabla c\right)\in L^{\tilde{p}}\left(\Omega_{T}\right)\,$,
@@
$\partial_{t}(h,d,m,c)\,,\,\Delta(h,d,m,c)\in[C^{0}\left(\Omega_{T}\right)]^{4}\,,\,\nabla\cdot\left(\chi(m,c)\nabla c\right)\in C^{0}\left(\Omega_{T}\right)\,.$
Moreover, there exists a constant $\overline{C}>0$, depending on $\Omega,h_{in},d_{in},m_{in},c_{in}$ and on all the system parameters, such that, the unique global solution of the system (4.1) satisfies:
$\sup_{t\geq 0}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[W^{1,\infty}(\Omega)]^{4}}\leq\overline{C}\,$, (4.6)
@@
The proof of Theorem 5 is divided into three parts. In Subsection 4.1 we prove the global existence of solutions by means of a Leray–Schauder fixed point argument. In Subsection 4.2 we establish uniqueness, while in Subsection 4.3 we derive the uniform-in-time bounds (4.6)–(4.7). The argument follows the approach developed in [5, 12], adapted here to the specific structure of the reaction-diffusion system (4.1).
@@
Let $1<p<\infty$, $0\leq t_{0}<T$, and $\Omega$ be a bounded domain in $\mathbb{R}^{n}$. Then there exists a constant $C_{T}>0$, depending on $T-t_{0}$, $\Omega$, $p$, and $n$, such that for all $u\in W^{1,2}_{p}(\Omega_{t_{0},T})$ the following estimates hold:
@@ 4.1.1 Construction of the map $S$
We begin by constructing a suitable solution operator $S$, which will be used to apply the Leray–Schauder fixed point theorem.
$S:[C^{0,1}]^{4}\times[0,1]\longrightarrow[C^{0,1}]^{4},$ (4.10)
as follows.
Given $(h,d,m,c)\in[C^{0,1}]^{4}$ and $\lambda\in[0,1]$, we define
@@
For a given $d\in C^{0,1}$, we consider the problem
[…]
where $d_{+}=\max(0,d)$.
Once $c^{\ast}$ is determined, for given $m,c\in C^{0,1}$ we solve
[…]
where $m_{+}=\max(0,m)$ and $c_{+}=\max(0,c)$.
Next, for given $h,d\in C^{0,1}$, we solve
[…]
@@
Finally, we consider
$\begin{cases}\partial_{t}h^{\ast}=\Delta h^{\ast}-h^{\ast}-A(m_{+},c_{+})h^{\ast}+\sigma(1-\overline{(h+d)}_{+})+\rho d^{\ast}(1-\overline{(h+d)}_{+})&(t,x)\in\Omega_{T},\\ \mathbf{n}\cdot\bm{\nabla}h^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\,,\\ h^{\ast}(x,0)=h_{in}(x)&x\in\Omega\,.\end{cases}$ (4.15)
@@
Moreover, from the definition of $S$ given by (4.12)–(4.15) and the maximal regularity theory for parabolic equations with Neumann boundary conditions (see [5, 21]), we obtain
@@
where $K^{\ast}$ depends on $T$, $\Omega$, $n$, $p$, the initial data, the parameters of the system, and the $C^{0,1}$ norms of $h$, $d$, $m$ and $c$.
Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (Lemma 6), and therefore
@@ 4.1.2 Compactness and continuity of $S$
To apply the Leray–Schauder fixed point theorem, we prove that the map $S$ is compact and continuous in $C^{0,1}]^{4}$.
We proceed in two steps. In Lemma 7, we show that $S$ maps bounded sets of $C^{0,1}]^{4}\times[0,1]$ into precompact sets of $C^{0,1}]^{4}$. In Lemma 8, we prove that $S$ is continuous.
@@
We start by estimating the difference $c^{\ast}_{1}-c^{\ast}_{2}$. From (4.12), using maximal regularity (see [21]) together with the embedding properties of $W^{1,2}_{p}$ (Lemma 6) for $p>n+2$, we obtain
@@
Applying the same arguments used for $c_{1}^{\ast}-c_{2}^{\ast}$, we deduce
@@
$\left|\rho d_{1+}\overline{(h_{1}+d_{1})}_{+}-\rho d_{2+}\overline{(h_{2}+d_{2})}_{+}\right|\leq\\ \leq\rho d_{2+}\left(|h_{1+}-h_{1+}|+|d_{1+}-d_{2+}|\right)+\rho(h_{1+}+d_{1+})|d_{1+}-d_{2+}|\,$, (4.20)
$\left|A(m_{1+},c_{1+})h_{1}^{\ast}-A(m_{2+},c_{2+})h_{2}^{\ast}\right|\leq\alpha(1+c_{\epsilon})\left(h_{2}^{\ast}|m_{1+}-m_{2+}|+m_{1+}|h_{1+}-h_{2+}|\right)+\\ +\alpha\,m_{1+}\,|h_{2}^{\ast}|\,|1-c_{\epsilon}||c_{1+}-c_{2+}|\,$, (4.21)
$|\overline{(h_{1}+d_{1})}_{+})-\overline{(h_{2}+d_{2})}_{+}|\leq(|h_{1+}-h_{2+}|+|d_{1+}-d_{2+}|)\,$, (4.22)
$\left|\rho d_{1}^{\ast}(1-\overline{(h_{1}+d_{1})}_{+})-\rho d_{2}^{\ast}(1-\overline{(h_{2}+d_{2})}_{+})\right|\leq\\ \leq\rho d_{2}^{\ast}\left(|h_{1+}-h_{2+}|+|d_{1+}-d_{2+}|\right)+\rho(1+h_{1+}+d_{1+}))|d_{1}^{\ast}-d_{2}^{\ast}|\,.$ (4.23)
Using (4.19)–(4.23) together with the same arguments as before, we obtain
@@ 4.1.3 The set $\Phi$
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m_{+},c_{+})h+\lambda\sigma(1-\overline{(h+d)}_{+})+\rho d(1-\overline{(h+d)}_{+}),&(t,x)\in\Omega_{T},\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c_{+})h+\rho\lambda d_{+}\overline{(h+d)}_{+},&(t,x)\in\Omega_{T},\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m_{+},c_{+})\nabla c\right)+\nu(\lambda d_{+}-m),&(t,x)\in\Omega_{T},\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d_{+}\,,&(t,x)\in\Omega_{T},\end{cases}$ (4.27)
with boundary conditions
$\mathbf{n}\cdot\bm{\nabla}h(t,s)=\mathbf{n}\cdot\bm{\nabla}d(t,s)=\mathbf{n}\cdot\bm{\nabla}m(t,s)=\mathbf{n}\cdot\bm{\nabla}c(t,s)=0\,\,\quad(t,s)\in(0,T)\times\partial\Omega\,$ (4.28)
and initial conditions
[…]
@@
[…]

### 03 Apr 2026 v3

@@ Abstract
We introduce a spatially extended mathematical model for Duchenne muscular dystrophy based on a damage-driven paradigm, in which immune recruitment is triggered by tissue injury. The model is formulated as a reaction–diffusion–chemotaxis system describing the interaction between healthy tissue, damaged fibers, immune cells and inflammatory signals. We establish the global well-posedness of the system and investigate the early-stage dynamics through linearization around the healthy equilibrium. Our analysis shows that diffusion does not induce Turing instabilities, so that spatial heterogeneity cannot arise from diffusion-driven mechanisms. Instead, disease progression occurs through invasion processes. We derive explicit conditions for the onset of invasion, interpreted as an effective damage reproduction threshold and characterize the minimal propagation speed of pathological fronts, showing that the dynamics is governed by a pulled-front mechanism. Numerical simulations support the analytical results and confirm the transition between decay and invasion. These results provide a mathematical framework for early-stage disease progression and indicate that spatial spreading arise from the expansion of localized damage rather than from intrinsic pattern-forming mechanisms.
@@ 3.2.2 Pathological equilibria
4 Well-posedness of the reaction-diffusion-chemotaxis system
In this section, we establish the global well-posedness of the reaction–diffusion–chemotaxis system (2.3). In particular, we prove the existence, uniqueness and boundedness of solutions, together with the preservation of a biologically admissible domain, ensuring that all components remain non-negative and that the saturation constraint $h+d\leq 1$ is satisfied for all times. The analysis relies on a Leray–Schauder fixed point argument, combined with a priori estimates and parabolic regularity techniques. These results provide the mathematical framework required for the subsequent investigation of the dynamical properties of the model.
@@ 4 Well-posedness of the reaction-diffusion-chemotaxis system
$\left\{\begin{array}[]{ll}\displaystyle\partial_{t}h=\Delta h+(\sigma+\rho d)(1-h-d)-h-A(m,c)h\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}d=D_{d}\Delta d+A(m,c)h-\delta d-\rho d(1-h-d)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(d-m)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}c=D_{c}\Delta c+rd-\mu c,&(t,x)\in\Omega_{T}\end{array}\right$. (4.1)
with:
$\chi(m,c)=\chi_{0}\dfrac{m\,c}{\kappa+c},\quad A(m,c)=\alpha\dfrac{c+c_{\epsilon}}{1+c}\,m\,$ (4.2)
and we denote by $\Omega_{T}=(0,T)\times\Omega$, where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ $(n\in\mathbb{N},\ n\geq 1)$ with smooth, specified later, boundary $\partial\Omega$. We write $\mathbf{n}=\mathbf{n}(s)$ for the outward normal to $\partial\Omega$ at a point $s$ of the boundary and impose the following Neumann boundary condition on $\partial\Omega$:
$\begin{cases}\mathbf{n}(s)\cdot\bm{\nabla}h(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}d(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}m(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}c(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\end{cases}$ (4.3)
@@
$\begin{cases}h(t=0,x)=h_{in}(x)\,,&\qquad x\in\Omega\cr d(t=0,x)=d_{in}(x)\,,&\qquad x\in\Omega\cr m(t=0,x)=m_{in}(x)\,,&\qquad x\in\Omega\cr c(t=0,x)=c_{in}(x)\,,&\qquad x\in\Omega\cr\end{cases}$ (4.4)
and we assume that the following condition holds:
@@
Moreover, we recall that all the system parameters are positive. We introduce the functional framework used throughout the section (see [15, 22]).
@@
We write $W^{r,p}(\Omega)$ for the Sobolev spaces of order $r$ with integrability $p$. For the parabolic setting, we consider the space:
@@
endowed with the norm:
@@
We also denote by $C^{k}(\overline{\Omega})$ the space of functions with continuous derivatives up to order $k$ and by $C^{0}(\Omega_{T})$ the space of continuous functions in $\Omega_{T}$.
We finally introduce the space $C^{0,1}=C^{0,1}([0,T]\times\overline{\Omega})$ of functions that are continuous together with their spatial gradients, endowed with the norm:
@@
$\partial_{t}(h,d,m,c)\,,\,\,\Delta(h,d,m,c)\,\in[L^{\tilde{p}}\left(\Omega_{T}\right)]^{4}\,\,,\,\,\quad\quad\nabla\cdot\left(\chi(m,c)\nabla c\right)\in L^{\tilde{p}}\left(\Omega_{T}\right)\,$
@@
$\partial_{t}(h,d,m,c)\,,\,\Delta(h,d,m,c)\in[C^{0}\left(\Omega_{T}\right)]^{4}\,,\,\,\quad\quad\nabla\cdot\left(\chi(m,c)\nabla c\right)\in C^{0}\left(\Omega_{T}\right)\,.$
Moreover, there exists a constant $\overline{C}>0$, depending on $\Omega,h_{in},d_{in},m_{in},c_{in}$ and on all the system parameters, such that the unique global solution of the system (4.1) satisfies:
$\sup_{t\geq 0}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[W^{1,\infty}(\Omega)]^{4}}\leq\overline{C}\,$ (4.6)
@@
The proof of Theorem 5 is divided into three parts. In section 4.1 we prove the global existence of solutions by means of a Leray–Schauder fixed point argument. In section 4.2 we establish uniqueness, while in section 4.3 we derive the uniform-in-time bounds (4.6)–(4.7). The argument builds on the approach developed in [6, 13], suitably adapted to the specific nonlinear and coupled structure of the reaction–diffusion system (4.1).
@@
Let $1<p<\infty$, $0\leq t_{0}<T$ and $\Omega$ be a bounded domain in $\mathbb{R}^{n}$. Then there exists a constant $C_{T}>0$, depending on $T-t_{0}$, $\Omega$, $p$ and $n$, such that for all $u\in W^{1,2}_{p}(\Omega_{t_{0},T})$ the following estimates hold:
@@ 4.1.1 Construction of the map $S$
We begin by constructing a suitable solution operator $S$, which will be used to apply the Leray–Schauder fixed point theorem, as follows:
$S:[C^{0,1}]^{4}\times[0,1]\longrightarrow[C^{0,1}]^{4}\,.$ (4.10)
Given $(h,d,m,c)\in[C^{0,1}]^{4}$ and $\lambda\in[0,1]$, we define:
@@
For a given $d\in C^{0,1}$, we consider the problem:
$\begin{cases}\partial_{t}c^{\ast}=D_{c}\Delta c^{\ast}-\mu c^{\ast}+rd_{+}&\quad(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}c^{\ast}=0&\quad(t,s)\in(0,T)\times\partial\Omega\\ c^{\ast}(t=0,x)=c_{in}(x)&\quad x\in\Omega\end{cases}$ (4.12)
where $d_{+}=\max(0,d)$. Once $c^{\ast}$ is determined, for given $m,c\in C^{0,1}$ we solve:
$\begin{cases}\partial_{t}m^{\ast}=D_{m}\Delta m^{\ast}-\nabla\cdot\left(\chi(m_{+},c_{+})\nabla c^{\ast}\right)+\nu(d_{+}-m^{\ast})&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}m^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ m^{\ast}(x,0)=m_{in}(x)&x\in\Omega\end{cases}$ (4.13)
where $m_{+}=\max(0,m)$ and $c_{+}=\max(0,c)$. Next, for given $h,d\in C^{0,1}$, we solve:
$\begin{cases}\partial_{t}d^{\ast}=D_{d}\Delta d^{\ast}-\delta d^{\ast}-\rho d^{\ast}+\rho d_{+}\overline{(h+d)}_{+}+A(m^{\ast},c_{+})h_{+}&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}d^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ d^{\ast}(x,0)=d_{in}(x)&x\in\Omega\end{cases}$ (4.14)
@@
Finally, we consider:
$\begin{cases}\partial_{t}h^{\ast}=\Delta h^{\ast}-h^{\ast}-A(m_{+},c_{+})h^{\ast}+\sigma(1-\overline{(h+d)}_{+})\\ \phantom{aaaa}+\rho d^{\ast}(1-\overline{(h+d)}_{+})&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}h^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ h^{\ast}(x,0)=h_{in}(x)&x\in\Omega\end{cases}$ (4.15)
@@
Moreover, from the definition of $S$ given by (4.12)–(4.15) and the maximal regularity theory for parabolic equations with Neumann boundary conditions (see [6, 22]), we obtain:
@@
where $K^{\ast}$ depends on $T$, $\Omega$, $n$, $p$, the initial data, the parameters of the system and the $C^{0,1}$ norms of $h$, $d$, $m$ and $c$.
Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (Lemma 6) and therefore:
@@ 4.1.2 Compactness and continuity of $S$
To apply the Leray–Schauder fixed point theorem, we show that the map $S$ is compact and continuous in $C^{0,1}]^{4}$. This is achieved by establishing precompactness of $S$ on bounded sets and then proving continuity in the following lemmas.
@@
We start by estimating the difference $c^{\ast}_{1}-c^{\ast}_{2}$. From (4.12), using maximal regularity (see [22]) together with the embedding properties of $W^{1,2}_{p}$ (Lemma 6) for $p>n+2$, we obtain:
@@
Applying the same arguments used for $c_{1}^{\ast}-c_{2}^{\ast}$, we deduce:
@@
$\left|\rho d_{1+}\overline{(h_{1}+d_{1})}_{+}-\rho d_{2+}\overline{(h_{2}+d_{2})}_{+}\right|\leq\\ \leq\rho d_{2+}\left(|h_{1+}-h_{2+}|+|d_{1+}-d_{2+}|\right)+\rho(h_{1+}+d_{1+})|d_{1+}-d_{2+}|\,$, (4.20)
$\left|A(m_{1+},c_{1+})h_{1}^{\ast}-A(m_{2+},c_{2+})h_{2}^{\ast}\right|\leq\alpha(1+c_{\epsilon})\left(|h_{2}^{\ast}||m_{1+}-m_{2+}|+m_{1+}|h_{1}^{\ast}-h_{2}^{\ast}|\right)+\\ +\alpha\,m_{1+}\,|h_{2}^{\ast}|\,|1-c_{\epsilon}||c_{1+}-c_{2+}|\,$, (4.21)
$|\overline{(h_{1}+d_{1})}_{+})-\overline{(h_{2}+d_{2})}_{+}|\leq(|h_{1}-h_{2}|+|d_{1}-d_{2}|)\,$, (4.22)
$\left|\rho d_{1}^{\ast}(1-\overline{(h_{1}+d_{1})}_{+})-\rho d_{2}^{\ast}(1-\overline{(h_{2}+d_{2})}_{+})\right|\leq\\ \leq\rho|d_{2}^{\ast}|\left(|h_{1}-h_{2}|+|d_{1}-d_{2}|\right)+\rho(1+h_{1+}+d_{1+}))|d_{1}^{\ast}-d_{2}^{\ast}|\,.$ (4.23)
Using (4.19)–(4.23) together with the same arguments as before, we obtain:
@@ 4.1.3 The set $\Phi$
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m_{+},c_{+})h+\lambda\sigma(1-\overline{(h+d)}_{+})+\rho d(1-\overline{(h+d)}_{+}),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c_{+})h+\rho\lambda d_{+}\overline{(h+d)}_{+},&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m_{+},c_{+})\nabla c\right)+\nu(\lambda d_{+}-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d_{+}\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.27)
with boundary conditions:
$\mathbf{n}\cdot\bm{\nabla}h(t,s)=\mathbf{n}\cdot\bm{\nabla}d(t,s)=\mathbf{n}\cdot\bm{\nabla}m(t,s)=\mathbf{n}\cdot\bm{\nabla}c(t,s)=0\,\,\quad(t,s)\in(0,T)\times\partial\Omega\$(4.28)
and initial conditions:
$\begin{cases}h(0,x)=\lambda h_{in}(x)&x\in\Omega\\ d(0,x)=\lambda d_{in}(x)&x\in\Omega\\ m(0,x)=\lambda m_{in}(x)&x\in\Omega\\ c(0,x)=\lambda c_{in}(x)&x\in\Omega\end{cases}$ (4.29)
@@
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m,c)h+\lambda\sigma(1-\overline{(h+d)}_{+})+\rho d(1-\overline{(h+d)}_{+}),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c)h+\rho\lambda d\overline{(h+d)}_{+},&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(\lambda d-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.30)